# Universal Approximation Theory: Foundations for Parallelism in Neural Networks

### v4

@@ Appendix B The Degree of Freedom of Different Layers Transformer
According to UAT2LLMs, the UAT format of multi-layer Transformer can be written as:
@@
where $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by a $j$-layer of UAT with $\mathbf{x}_{0}$ as input. All $\mathbf{W}^{\prime}_{i,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input. Because they are affected by the weight of MHA in each layer. Why does the MHA operation affect the weight parameters in a linear manner? According to Eq. (12) and Eq. (13), this is because $\mathbf{W}^{\prime}_{1,2}=\underline{\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ represents the MHA parameters, which are influenced by the input, while $\mathbf{W}^{\prime}_{1,2}$ is a defined parameter. As a result, the parameter $\mathbf{W}^{\prime}_{1,2}$ in Eq. (13) becomes input-dependent, and this influence occurs through a linear transformation. Similarly, in a multi-layer Transformer, multiple layers of weight parameters are derived through similar transformations.
Therefore, it is easy to deduce that the degrees of freedom for Transformers with 1, 2, 3, and 6 layers are 2, 6, 9, and 18, respectively. For Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24, the degrees of freedom are shown in Table 1. At first glance, it seems that the degrees of freedom for Para-F-1-24, Para-F-2-24, Para-F-3-24, and Para-F-6-24 all exceed that of Para-F-6-6. Why then does only Para-F-6-24 consistently outperform Para-F-6-6?
@@
The reason lies in the nature of the degrees of freedom. Taking $\mathbf{W}^{\prime}_{2,2}$ and $\mathbf{b}^{\prime}_{2,2}$ from Eq. (15) as examples, $\mathbf{W}^{\prime}_{2,2}=\underline{\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}}$, where $\mathbf{W}^{\prime}_{1,1}$ is an input-dependent parameter from MHA, and $\mathbf{W}^{\prime}_{2,2}$ is a defined parameter, becoming fixed post-training. Thus, the degrees of freedom for $\mathbf{W}^{\prime}_{2,2}$ in Eq. (15) are imparted through a linear transformation with the MHA parameter $\mathbf{W}^{\prime}_{1,1}$. Additionally, $\mathbf{b}^{\prime}_{2,2}=\underline{(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})}$. Hence, $\mathbf{b}^{\prime}_{2,2}$ in Eq. (15) can be understood as being approximated by a two-layer UAT, showcasing stronger dynamic transformation capabilities based on inputs. Due to UAT’s powerful capacity, $\mathbf{b}^{\prime}_{2,2}$ can dynamically fit specific functions based on inputs.
In a single-layer Transformer, as seen in Eq. (12), $\mathbf{b}^{\prime}_{1,3}$ and $\mathbf{b}^{\prime}_{1,2}$ are defined parameters, so the number of bias terms approximated using UAT is zero. For a two-layer Transformer, as shown in Eq. (15), the number of bias terms approximated using UAT is two, specifically $\mathbf{b}^{\prime}_{2,1}$ and $\mathbf{b}^{\prime}_{2,2}$, with the UAT layers approximating these bias terms being 1 and 2, respectively. According to Eq.(16), in a Transformer with $i+1$ layers, $\mathbf{b}_{i+1,1}$ is approximated by a single-layer UAT, while $\mathbf{b}^{\prime}_{j,2}$ is approximated by a UAT of $j$ layers for $j>1$. It is easy to deduce that for a three-layer Transformer, the number of bias terms approximated using UAT is three, with the UAT layers being 1, 2, and 3. For a six-layer Transformer, the number of bias terms approximated using UAT is six, with the UAT layers being 1, 2, 3, 4, 5, and 6. The UAT layers of bias terms for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are shown in Table 2.
@@
Table 2: The UAT layer count corresponding to the bias terms in Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 is calculated by summing the UAT layers associated with each bias term within a serial Transformer. Since each serial Transformer has multiple bias terms, with each term corresponding to different UAT layer counts, we represent the UAT layers by summing the UAT layers for each bias term. For example, ”1+2+3” indicates that three bias terms are approximated using UAT, with 1, 2, and 3 layers respectively. Additionally, since multiple serial Transformers are used in parallel, we multiply the sum by the number of parallel Transformers.

### v5

@@ Appendix B The Degree of Freedom of Different Layers Transformer
According to DUAT2LLMs, the DUAT format of a multi-layer ViT can be written as:
@@
When $i=0$, the parameters are defined as follows: $\mathbf{W}^{\prime}_{1,1}=\mathbf{W}^{\prime}_{1,1}$, $\mathbf{b}^{\prime}_{1,3}=\mathbf{b}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,3}=\mathbf{W}^{\prime}_{1,3}$, $\mathbf{W}^{\prime}_{1,2}=\mathbf{W}^{\prime}_{1,2}\mathbf{W}^{\prime}_{1,1}$, and $\mathbf{b}^{\prime}_{1,2}=\mathbf{b}^{\prime}_{1,2}$. For cases where $i\geq 1$, we define the parameters iteratively. For each $j=1,2,\ldots,i$, the updates are as follows: $\mathbf{W}^{\prime}_{j+1,1}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{W}_{j,1}$, $\mathbf{b}^{\prime}_{j+1,3}=\mathbf{W}^{\prime}_{j+1,1}\mathbf{b}_{j,1}+\mathbf{b}^{\prime}_{j+1,3}$, $\mathbf{W}^{\prime}_{j+1,2}=\mathbf{W}^{\prime}_{j+1,2}\mathbf{W}_{j,1}$, and $\mathbf{W}^{\prime}_{j,3}=\mathbf{W}^{\prime}_{i+1,1}\mathbf{W}^{\prime}_{i,1}\cdots\mathbf{W}^{\prime}_{j,3}$. Additionally, for $j=1,2,\ldots,i+1$, the bias terms $\mathbf{b}^{\prime}_{j,2}$ are updated according to:
$\mathbf{b}^{\prime}_{j,2}=$ $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})$ (12)
$+$ $\mathbf{W}^{\prime}_{j,2}\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2}).$
The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by a $j-1$-layer of DUAT with $\mathbf{x}_{0}$ as input. The parameters $\mathbf{W}^{\prime}_{j,1}$ and $\mathbf{W}^{\prime}_{j,2}$ for $i+1\geq j\geq 1$, and $\mathbf{W}^{\prime}_{j,3}$ for $i+1>j>1$ in layer $i$, all adapt dynamically based on the input. Because they are affected by the weight of MHA in each layer.
Therefore, it is easy to deduce that the degrees of freedom for Transformers with 1, 2, 3, and 6 layers are 2, 6, 9, and 15, respectively. For Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24, the degrees of freedom are shown in Table 1. At first glance, it seems that the degrees of freedom for Para-F-1-24, Para-F-2-24, Para-F-3-24, and Para-F-6-24 all exceed that of Para-F-6-6. Why then does only Para-F-6-24 consistently outperform Para-F-6-6?
@@
The reason lies in the nature of the degrees of freedom. Taking $\mathbf{W}^{\prime}_{2,2}$ and $\mathbf{b}^{\prime}_{2,2}$ as examples, $\mathbf{W}^{\prime}_{2,2}=\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,1}$, where $\mathbf{W}^{\prime}_{1,1}$ is an input-dependent parameter from MHA, and $\mathbf{W}^{\prime}_{2,2}$ is a defined parameter, becoming fixed post-training. Thus, the degrees of freedom for $\mathbf{W}^{\prime}_{2,2}$ is imparted through a linear transformation with the MHA parameter $\mathbf{W}^{\prime}_{1,1}$. Additionally, $\mathbf{b}^{\prime}_{2,2}=(\mathbf{W}^{\prime}_{2,2}\mathbf{b}^{\prime}_{1,3}+\mathbf{b}^{\prime}_{2,2})+\mathbf{W}^{\prime}_{2,2}\mathbf{W}^{\prime}_{1,3}\sigma(\mathbf{W}^{\prime}_{1,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{1,2})$. Hence, $\mathbf{b}^{\prime}_{2,2}$ can be understood as being approximated by a two-layer DUAT, showcasing stronger dynamic transformation capabilities based on inputs. Due to DUAT’s powerful capacity, $\mathbf{b}^{\prime}_{2,2}$ can dynamically fit specific functions based on inputs.
In a single-layer Transformer, as seen in Eq. equation 10, $\mathbf{b}^{\prime}_{1,3}$ and $\mathbf{b}^{\prime}_{1,2}$ are defined parameters, so the number of bias terms approximated using UAT is zero. For a two-layer Transformer, the number of bias terms approximated using DUAT is one, specifically $\mathbf{b}^{\prime}_{2,2}$ with the DUAT layers approximating this bias terms being 1. According to Eq.equation 11, in a Transformer with $i+1$ layers, $\mathbf{b}_{i+1,1}$ is approximated by a single-layer UAT, while $\mathbf{b}^{\prime}_{j,2}$ is approximated by a UAT of $j$ layers for $j>1$. It is easy to deduce that for a three-layer Transformer, the number of bias terms approximated using UAT is three, with the UAT layers being 1, 2, and 3. For a six-layer Transformer, the number of bias terms approximated using DUAT is six, with the DUAT layers being 1, 2, 3, 4, 5, and 6. The DUAT layers of bias terms for Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 are shown in Table 2.
@@
Table 2: The DUAT layer count corresponding to the bias terms in Para-F-1-24, Para-F-6-6, Para-F-2-24, Para-F-3-24, and Para-F-6-24 is calculated by summing the UAT layers associated with each bias term within a serial Transformer. Since each serial Transformer has multiple bias terms, with each term corresponding to different UAT layer counts, we represent the UAT layers by summing the UAT layers for each bias term. For example, ”1+2+3” indicates that three bias terms are approximated using UAT, with 1, 2, and 3 layers respectively. Additionally, since multiple serial Transformers are used in parallel, we multiply the sum by the number of parallel Transformers.